OpenStudy (anonymous):

y = 2x + 4 2x - y = 6 Solve the system of equations using substitution. A) (2,8) B) (2,-8) C) (10,24) D) no solution

OpenStudy (anonymous):

@swagmaster47

OpenStudy (anonymous):

So we know that y=2x+4 so lets replace y with 2x+4 in the second equation 2x-y=6 Becomes ... 2x-(2x+4)=6 Solve for x

OpenStudy (anonymous):

I think I got it it might be uh A)2,8 if I got it wrong thanks still for helping!

OpenStudy (anonymous):

Following up from what Heart_Broken_Kid said, continue solving for x and see what happens.

OpenStudy (anonymous):

Hmmm lets solve for x 2x-(2x+4)=6 2x-2x-4=6 0x-4=6 -4=6 -4 CAN NOT EQUAL 6 THIS IS FALSE THEREFORE THERE IS NO SOLUTION